OpenStudy (anonymous):

Assume a 125 pound person burns 150 calories after 15 minutes of jogging and 200 calories after 20 minutes of jogging. Part 1 (2 points): Use calories as the y-coordinate and minutes as the x-coordinate. After writing two ordered pairs, find the slope. What does the slope represent in terms of the information provided? Part 2 (2 points): Write an equation, in slope intercept form, to represent this data. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you need to find the equation of the line that goes through (15, 150) and (20,200)

OpenStudy (anonymous):

50/5?

jimthompson5910 (jim_thompson5910):

which reduces to what

OpenStudy (anonymous):

10?

jimthompson5910 (jim_thompson5910):

that's the slope

jimthompson5910 (jim_thompson5910):

so y = mx + b y = 10x + b now plug in either point, then solve for b

OpenStudy (anonymous):

150=10(15)+b?

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (anonymous):

150=150+b? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

then what

OpenStudy (anonymous):

um im not sure

jimthompson5910 (jim_thompson5910):

150+b is the same as b+150

jimthompson5910 (jim_thompson5910):

150 = b+150 .... b = ??

OpenStudy (anonymous):

150?

jimthompson5910 (jim_thompson5910):

150+SOMETHING = 150 what must this SOMETHING be?

jimthompson5910 (jim_thompson5910):

to find out, you can subtract 150 from both sides.

OpenStudy (anonymous):

so wouldnt it be 0?

jimthompson5910 (jim_thompson5910):

b = 0, yep

jimthompson5910 (jim_thompson5910):

m = 10 b = 0 y = mx+b y = 10x + 0 y = 10x

OpenStudy (anonymous):

How many calories will this person burn after 35 minutes of jogging? Using complete sentences, explain how the equation, slope, or graph can help to predict calories burned. what about this?

jimthompson5910 (jim_thompson5910):

plug x = 35 into y = 10x and you tell me

OpenStudy (anonymous):

350?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

so you can see how the equation is used to predict how many calories are burned

jimthompson5910 (jim_thompson5910):

the graph is just a visual way of doing the same thing pretty much

OpenStudy (anonymous):

so the slope represents the number of calories burned?

jimthompson5910 (jim_thompson5910):

per minute

jimthompson5910 (jim_thompson5910):

so if you workout for one more additional minute, you burn 10 more calories

OpenStudy (anonymous):

ohh okay i get it now thank you so much!(:

jimthompson5910 (jim_thompson5910):

yw